lisa traveled to her cabin on the lake and back. the trip there took five hours and the trip back took six hours. she averaged 45 kilometers per hour on the return trip. find the average speed of the trip there

Answer by lwsshak3(11628)

lisa traveled to her cabin on the lake and back. the trip there took five hours and the trip back took six hours. she averaged 45 kilometers per hour on the return trip. find the average speed of the trip there.
***
Ave speed=distance/travel time
distance=2*45*6=540 km (lake and back)
travel time=5+6=11 hrs
average speed of the trip=540/11≈49.09 km/hr